### v1

# Quasi-Deterministic Modeling of Sub-THz Band Access Channels in Street Canyon Environments

Abstract
Sub-terahertz (sub-THz) frequencies (100–300 GHz) are expected to play a key role in beyond-5G and 6G mobile networks. However, their quasi-optical propagation characteristics require new channel models beyond sub-100 GHz extrapolations. This paper presents an extensive double-directional (D-D) channel measurement campaign conducted in an outdoor street-canyon environment at 154 GHz and 300 GHz under both line-of-sight (LoS) and non-line-of-sight (NLoS) conditions using an in-house-developed channel sounder. Based on these measurements, clustering with merged datasets across the two frequencies enables comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (QD) channel model is then proposed, combining deterministic components, such as LoS and single-bounce reflections from side walls, with random components. Large-scale parameters (path loss, delay spread, angular spread, and Rician K-factor) are also evaluated. These results provide valuable insights into sub-THz propagation in urban street canyons and contribute toward the development of accurate, channel models for future 6G systems.
Index Terms:
Sub-terahertz, dual-band channel sounding, outdoor multipath propagation, Quasi-deterministic channel model, street canyon, outdoor hot spot, multipath interactionI Introduction
Sixth-generation (6G) mobile communications are expected to achieve ultra-high data rates exceeding Gbps and ultra-low latency on the order of s, far surpassing the capabilities of current networks. To meet these stringent requirements, research has increasingly focused on the low-terahertz (low-THz) or sub-THz spectrum (– GHz), which offers contiguous channel bandwidths of several tens of GHz. A significant milestone was reached at the World Radiocommunication Conference 2023 (WRC-23), where it was agreed to consider allocating the – GHz, – GHz, – GHz, – GHz, and – GHz bands for future International Mobile Telecommunications (IMT) as part of the preliminary agenda for WRC-31 [1]. This regulatory progress, together with growing research in channel measurements, propagation modeling, and early standardization discussions in organizations such as 3GPP [2] and ETSI [3], has further strengthened expectations that sub-THz frequencies will become a cornerstone of mobile networks beyond 2030.
Although THz frequencies have so far been applied mainly to point-to-point (P2P) fixed links and backhaul communications, their extension to short-range mobile access is increasingly recognized as promising. According to a 2024 ETSI THz Industry Specification Group (ISG) report [3], THz communications could enable a wide range of transformative use cases. However, realizing these applications requires addressing the quasi-optical propagation behavior of THz waves and mitigating their severe propagation losses. THz channels are strongly dominated by line-of-sight (LoS) propagation and experience heavy attenuation in non-line-of-sight (NLoS) conditions due to blockage by both static and dynamic objects. Reliable 100-Gbps mobile access therefore depends on highly directional links with a combined transmitter (Tx) and receiver (Rx) antenna gain exceeding 50 dBi, capable of exploiting both the LoS path and multipath components [4]. Yet, reflection losses, amplified by rough-surface scattering, limit multipath viability as the distance between antennas and reflectors increases [5]. Densifying base station deployments can mitigate coverage gaps but is economically unfeasible. Instead, exploiting multipath through multi-beam management and passive/intelligent reflecting surfaces is essential, though intelligent reconfigurable surface (IRS) technologies still face challenges in control, placement, and cost. A fundamental understanding of sub-THz multipath channels is therefore critical.
Sub-THz hot spots in street environments provide ultra-high-capacity access in dense urban areas, as illustrated in Fig. 1, enabling transformative applications such as immersive augmented reality (AR), virtual reality (VR), mixed reality (MR), connected mobility, and real-time edge computing. Several measurement campaigns have recently investigated sub-THz propagation in urban street canyon environments to support future 6G channel modeling. Angle-resolved measurements at and GHz in a UMi street canyon were conducted, highlighting the sparse multipath structure of THz channels [6]. Similar results were obtained in double-directional (D-D) measurement campaigns at 145 GHz, which revealed dominant LoS and facade reflections with limited angular spread [7]. At 140 and 220 GHz, LoS path loss was found to approximate free-space predictions, while NLoS conditions suffered significantly higher attenuation due to building obstructions. Multipath richness was still observed, with notable contributions from facades, vehicles, and street-level scatterers [8]. By contrast, 306–371 GHz measurements in a campus street environment revealed much sparser propagation: wall reflections incurred 5–30 dB additional loss, and both delay and angular spreads were substantially smaller than those reported in sub-100 GHz 3GPP street-canyon models. Collectively, these results highlight a clear trend: as frequency increases from 140 to 370 GHz, multipath richness diminishes and channels become increasingly sparse, underscoring the need for new channel models tailored to THz propagation [9]. Additional large-scale efforts by collaborative studies in ETSI and ITU-R have provided datasets [10, 11] that are now feeding into early sub-THz channel model development for 3GPP. Collectively, these studies demonstrate that street hot spots are crucial deployment scenarios for THz access, where LoS dominance and a few strong reflections can be effectively exploited for reliable high-capacity connectivity.
As explained above, recent sub-THz channel measurement campaigns have provided valuable insights into propagation in urban street environments; however, several important gaps remain. While path loss and other large-scale characteristics have been reported, detailed multipath analyses are largely missing, leaving open questions about the number and site-specific nature of multipath components that can be effectively exploited. Integrating multiple-input-multiple-output (MIMO) antennas into THz systems enables ultra-narrow directional beamforming to mitigate severe path loss, and multi-stream transmission via multi-beam MIMO can significantly enhance data rates. Consequently, understanding multipath cluster channel characteristics at THz frequencies is essential for accurately assessing the intrinsic channel capacity of an environment and expanding coverage with IRS. From a modeling perspective, current standardized frameworks such as 3GPP TR 38.901 [2] still rely on extrapolations from sub-100 GHz data and assume rich multipath scattering, which does not reflect the sparse structure of sub-THz channels. Consequently, channel model representations that accurately capture frequency-dependent cluster behavior and quasi-deterministic propagation mechanisms in street canyons are still lacking [12].
In this context, the present study reports an extensive D-D channel measurement campaign conducted in an outdoor street-canyon environment at 154 and GHz under both LoS and NLoS conditions using an in-house-developed channel sounder. Based on these measurements, clustering with merged datasets from the two frequency bands is performed, enabling comparative analyses that identify both common and distinct multipath clusters, as well as the frequency dependence of cluster-level characteristics. A quasi-deterministic (QD) channel model is then proposed, in which the deterministic components include the LoS path and single-bounce reflections from the two side walls, while the randomness of multipath components is modeled using an approach adapted from corridor scenarios [13] owing to their structural similarity. The deterministic components also incorporate randomness introduced by reflector’s surface conditions, represented by a two-state Markov process for presence and absence. Finally, the propagation channel is comprehensively characterized in terms of large-scale parameters (LSPs), including path loss (PL), delay spread (DS), azimuth angular spread (AS), and Rician K-factor (KF). The main contributions of this work are summarized as:
Extensive dual-band simultaneous D-D channel measurements at 154 and GHz in a street-canyon environment under both LoS and NLoS conditions using a consistent setup;
Dual-band joint signal processing, including noise filtering for enhanced multipath detection and clustering for comparative analysis across the two frequency bands;
Development of a QD channel model with deterministic components and random components modeled using a corridor-inspired approach; and
Comprehensive evaluation of large-scale parameters, namely PL, DS, AS, and Rician KF.
The remainder of this paper is organized into five sections. Section II outlines the measurement campaign, including scenario descriptions and a brief overview of the channel sounding setup. Section III details the post-processing procedures, such as dual-band joint noise filtering and clustering. Section IV examines the dominant propagation mechanisms by analyzing the trajectories of measured power spectra and validating them through ray-tracing. Section V presents the proposed QD channel model, while Section VI characterizes the channel in terms of path loss (PL) and large-scale parameters (LSPs). Finally, Section VII summarizes the key findings and highlights directions for future research.
II Measurement campaign
II-A Measurement Scenarios
The D-D channel measurements were conducted in a street canyon with a width of 20–25 m on the Niigata University campus. High-rise buildings (height m) line both sides of the street canyon, as illustrated in Fig. 2(a). In addition, there are large glass windows on the side walls. During the measurement, the Rx was placed in 30 distinct positions (Rx1–Rx30) at a height of m from the ground, emulating a mobile terminal. The Tx was fixed at a height of m above the ground and sequentially positioned at two locations at one end of the street canyon. One location provided a LoS scenario, with Tx1 corresponding to receiver positions Rx1–Rx23 (distance ranging from 10–80 m). The other location was positioned behind the corner of a building, blocking the LoS path to the receiver and creating an NLoS scenario, with Tx2 corresponding to receiver positions Rx24–Rx30 (distance ranging from 10–40 m).
Directional scanning measurement was conducted, and the setup is shown in Fig. 2(b). On both sides, dBi gain directional horn antennas with the narrow HPBWs of and were employed for both bands. The full azimuth (Az) sweep was conducted at both Rx and Tx, whereas the Rx and Tx antenna elevation (El) angles were fixed at and , respectively ( tilted). As a result, angular samples in the azimuth at both Tx and Rx were obtained (1,600 combinations). The antenna gains and radiation pattern effects were removed from the measured data in the post-processing.
II-B Channel Sounding
Measurements were conducted using a GHz dual-band channel sounder developed from our earlier work [14]. Since the 300-GHz system has already been described in detail in [14], this paper highlights the dual-band extensions and their distinctive features. The system parameters are summarized in Table I.
At the Tx side, two intermediate-frequency (IF) signals with a center frequency of 8 GHz are generated using two separate channels of an arbitrary waveform generator (AWG, M8195A, Keysight) operating at a sampling rate of 64 GSa/s. The signal generator (SG) produces a local oscillator (LO) signal at 24.33 GHz, which is fed into two frequency upconverters (WR6.5 CCU and WR3.4 CCU, VDI Corporation). These upconverters multiply the LO frequency by factors of 6 and 12, thereby converting the IF signals to the GHz and GHz bands, respectively. Each dual-band signal is then filtered by a band-pass filter (BPF) to suppress image frequencies, amplified by a power amplifier (PA), and transmitted via a horn antenna. At the Rx side, the signals are captured by horn antennas and downconverted using WR6.5 and WR3.4 CCD modules (VDI Corporation) for the 154 and GHz bands, respectively. The resulting IF signals are amplified by low-noise amplifiers (LNAs) and sampled with a high-speed digitizer (M8131A, Keysight) for storage and subsequent processing. Multi-tone sounding signals were employed, consisting of 2,560 tones over 4 GHz bandwidth at GHz and 5,120 tones over 8 GHz bandwidth at GHz. This configuration provided delay resolutions of 250 ps and 125 ps for the 154 and GHz bands, respectively, while maintaining a common delay span of 640 ns.
Parameters | Freq I | Freq II | ||||||
Center frequency | GHz | GHz | ||||||
Signal bandwidth | 4 GHz | 8 GHz | ||||||
No. Subcarriers, | ||||||||
FFT points, | ||||||||
Sampling rate |
Delay resolution | 250 ps | 125 ps | ||||||
Delay span | 640 ns | |||||||
Horn antenna |
Polarization | Vertical |
To compensate for the effects of system response, a back-to-back (B2B) calibration was performed using waveguide attenuators (ATT) prior to the campaign [14]. The effect of the antenna gain was removed through over-the-air (OTA) calibration by aligning the antennas, and the tilting effect was also compensated for by post-processing. Furthermore, considering the propagation loss associated with the distance between the Tx and Rx, the signal-to-noise ratio (SNR) of the received signal in each band is expected to degrade significantly. This limitation restricts the achievable measurement range. To mitigate this issue, the continuously received sounding symbols are coherently averaged over multiple symbol durations, thereby improving the SNR and extending the effective measurement range.
III Post-Processing
III-A Data format
The data acquired in the measurement is the angle-resolved channel transfer function (CTF), , where , , and are the frequency bin, and the pointing azimuth angles of departure (AoD) and arrival (AoA), respectively. The channel impulse response (CIR), , is obtained by taking the inverse Fourier transform of the measured CTF, where is the delay bin. The omnidirectional power delay spectrum (PDP), the azimuth delayed power spectra (ADPS), and the angular power spectra (APS) can be obtained by summing the relevant spectral bins as
where is the noise-filtered double-directional angular delay power spectrum (DDADPS) [14, 13], denotes the sample index of the subscripted domain, and , is the complementary to . The DDADPS is defined as the expected value of the square of the absolute CIR as
where represents the expectation operator.
III-B Noise filtering for dual-band datasets
The measured CIR sample set consists of both signal and noise components, namely , where and denote the signal and noise sample sets, respectively. Owing to the sparsity of the sub-THz channel, only a few samples correspond to actual signal components, whereas the majority represent noise. Effective removal of the noise component is therefore essential for reliable channel analysis. The threshold level for noise filtering can be determined from the estimated noise variance, since the noise samples follow a complex Gaussian distribution with zero mean and variance , implying that the noise power follows an exponential distribution [15, 16].
The noise variance can be estimated as
where
and denote the empirical cumulative distribution function obtained using all samples and the exponential distribution model fitted by the smallest samples in (ascending-sorted version of ), respectively. Then, the threshold level is calculated as
where is the target confidence level, defined as
and is a margin ensuring no false detection is expected when testing the samples times the total number . Here, was chosen.
The common threshold in (7), applied to all samples in a single dataset, generally yields satisfactory performance. However, sample sets at two frequencies, even when processed under identical noise-filtering conditions, often produce CIRs that do not align well due to frequency-dependent behavior or false detections. This suggests that noise filtering can be enhanced by cross-checking the two datasets, leveraging the fact that propagation mechanisms remain fundamentally consistent across frequencies. In this approach, the threshold for each frequency is adaptively adjusted along the delay axis by increasing or decreasing the value from the common threshold at delay bins where samples should be retained or removed, respectively. The detailed procedure is provided in Algorithm 1, where , denoting the PDP value of at the delay bin corresponding to , is compared with the common threshold of slightly reduced by , i.e., . If this value exceeds the reduced threshold, the individual threshold at is updated as to revive a miss-removed path; otherwise, the individual threshold at is updated as to eliminate a miss-detected path.
Fig. 3 shows an example of clustering (Rx19), where all CIR samples are plotted along the delay axis, with the decision margin empirically set to dB. The results demonstrate that multipath components are consistently detected across the two frequencies. It is noted that the resulting dynamic range (DR), defined as the difference between the peak power and the noise floor (), ranges from approximately to dB, while the spurious-free dynamic range (SFDR), defined as the difference between the peak power and the threshold level, ranges from to dB in the LoS scenario.
III-C Clustering
The multipath components (MPCs) were extracted using the Sub-grid CLEAN algorithm [16], which estimates the parameters of each MPC by sequentially subtracting a replica from the DDADPS in descending order of power. Each replica is generated using the antenna radiation pattern and the signal autocorrelation function. Clustering was applied to the extracted MPCs to group them into clusters based on similar angle and delay parameters. The clustering process was performed on the merged datasets of MPCs obtained at both frequencies under a consistent setup, in order to assign identical cluster indices across the two frequency sample sets [17]. To account for the increased propagation loss due to the frequency difference, the power of each GHz MPC was scaled up by dB prior to clustering. After clustering, the clustered MPCs were separated back into their respective frequency datasets, and the power of each GHz MPC was restored to its original value. This approach enabled the identification of both common clusters and uncommon (frequency-specific) clusters. The -PowerMeans clustering algorithm was employed, with the number of clusters, , determined manually by visual inspection to preserve the physical interpretability of the results. Fig. 4 presents an example clustering result for Rx19, showing four clusters that are common across both frequency bands.
IV Multipath Cluster Behavior
IV-A Dominant Scattering Process
Fig. 5(a) shows photos of the surrounding buildings located to the north and south. The building walls, composed of glass windows/doors and block facades, serve as the primary reflectors that generate the strong single-bounce (SB) paths observed in the measurement campaign. It is noted that the north wall is not perfectly parallel to the street, and there are no trees in front of the walls. Fig. 5(b) illustrates a simplified two-dimensional (2D) geometrical model of the LoS scenario, which is used for ray-tracing calculations. The material constants for both walls, primarily composed of concrete, are characterized by the following relative complex permittivities: at GHz and at GHz [18].
Figs. 6 depicts the spatial variation of the PDPs along the Rx trajectory for both scenarios. Specifically, Fig. 6(a) shows the relative PDPs, where the delay is normalized to the LoS delay, while Fig. 6(b) presents the normal PDPs as a function of the absolute delay. At both frequencies, the propagation characteristics exhibit pronounced sparsity and simplicity. In the LoS scenario, as shown in Fig. 6(a), two major scattering processes that vary consecutively along the Rx trajectory are identified as SB reflections from the north wall (NW) and south wall (SW), with their trajectories validated using the image method ray-tracing applied to the simplified model in Fig. 5(b). Aside from the three main components, such as the LoS path and the two SB reflections, only a few additional comonents are observed, mainly originating from double-bounce (DB) reflections and other unidentified interactions with random objects. This phenomenon become more pronounced at GHz. In contrast, in the NLoS scenario, as shown in Fig. 6(b), two dominant propagation mechanisms that evolve consecutively along the Rx trajectory are approximately identified as the SB reflection from the SW and the DB reflection resulting from the SW-NW interaction.
The behavior of the extracted MPCs is illustrated in Figs. 7(a), 7(b), and 7(c), which present the variations in MPC distribution across the delay, AoD, and AoA domains, respectively, at both frequencies. In the LoS scenario, the SB reflections from the NW and SW dominate, whereas DB reflections are rarely observed, a trend that was further confirmed by the clustering results.
IV-B Cluster Characteristics
Fig. 8 illustrates the cluster characteristics through the cumulative distribution function (CDF). Fig. 8(a) depicts the distribution of the number of clusters, showing that the GHz band exhibits slightly more clusters than the GHz band. This difference can be attributed to the higher interaction losses at GHz. It is also observed that the number of clusters in the NLoS scenario is slightly smaller than that in the LoS scenario.
Fig. 8(b) shows the path gain (PG) distribution of the clusters for each frequency band. The solid lines represent the LoS cluster PG across all Rx points in the LoS scenario, while the dashed lines denote the total PG of the NLoS clusters. For the LoS clusters, the average PG values are dB at GHz and dB at GHz, corresponding to a power difference of dB. This difference is dB smaller than the additional FSPL of dB, indicating that the GHz band experiences slightly greater degradation. However, for the NLoS clusters, the average PG values are dB at GHz and dB at GHz, yielding a power difference of dB, which is dB larger than the additional FSPL.
As explained above, the PGs of the NLoS clusters at GHz exhibit slightly greater degradation compared to those at GHz. To further examine this, Fig.8(c) presents a comparative analysis based on the relative powers of common clusters, which helps mitigate practical calibration issues such as inaccurate antenna gain. The results show that the statistical distribution of NLoS cluster power relative to the LoS cluster power is quite similar for both frequency bands. Moreover, the dashed black curve, representing the difference between the two frequencies, remains roughly balanced, indicating no significant disparity between the GHz and GHz bands, despite the wide distribution range from to dB. It is also observed, however, that degradation is more pronounced at GHz when the relative path gain difference is positive.
V Channel modeling
Based on the investigation of the propagation mechanisms, this section proposes a QD channel model inspired by those developed for corridor scenarios [13]. The discussion begins with the mathematical representation of the channel, followed by the modeling and generation of the deterministic and random components that constitute the channel. It should be noted that the channel model is developed only for the LoS scenario, due to the limited amount of available data.
V-A QD channel model
Due to the short wavelength of sub-THz waves, specular reflection dominates and can be effectively predicted using deterministic methods such as ray tracing. Building on this characteristic as the deterministic component, a QD channel model[5, 12] further accounts for non-negligible interactions arising from higher-order reflections and scattering caused by object shape and surface roughness, modeling them as random components. Namely, the propagation channel can be modeled as
where the deterministic components, , and random components, are separately represented.
As inferred in the previous section, the deterministic components are represented by the LoS path and the two dominant wall reflections from the NW and SW, and can be expressed as
where ‘’ and ‘’ denote the single-bounce reflections by NW and SW, respectively. denotes a random variable that captures the randomness of the path amplitude, including its presence or absence.
Freq. | Transision probability matrix |
GHz | , |
GHz | , |
V-B Deterministic components
When comparing the measurement results with ray-tracing simulations, it is often observed that certain reflected components appear randomly in the measurements and do not align well with the simulated results, thereby generating power differences, as illustrated Fig. 9. A plausible cause of this trend may be the inhomogeneity in the wall shape and surface smoothness, along with the presence of various glass components such as windows and doors.
Fig. 9(a) illustrates the distribution of the differences between the measured and calculated powers for the SB reflections from the NW and SW at both frequencies, where missing Rx positions indicate the absence of a signal. Fig.9(b) further presents the empirical CDFs together with their fitted models. To capture this probabilistic behavior induced by factors such as surface roughness, these models can be combined with a two-state Markov chain to represent signal presence or absence, as proposed in [13]. This property of the Markov chain is represented by the transition probability matrix . The modeling of path appearance is carried out separately for the SB reflections from the NW and SW at each frequency. Table II summarizes the estimated transition probability matrices for each reflector.
Freq. | 1/ | |||
GHz | ||||
GHz |
V-C Random components
As described above, non-negligible random interactions should be incorporated into the channel model for a more accurate representation. Fig. 10 shows the distributions of relative PG for the deterministic and random components, where the SB reflections from the NW and SW are treated as deterministic components and all other components are treated as random. Evidently, the random components make a significant contribution to the overall power. As explicated earlier, MPCs originating from non-predominant higher-order reflections and difficult-to-identify propagation paths are modeled as random components.
Following the procedure in [13], Fig. 11(a) presents the CDF of the interarrival times of these random components at both frequencies. The distribution of interarrival time is expected to follow an exponential distribution as
and has been fitted accordingly in Fig. 11(a). Here, represents the arrival time of the th MPCs, and is the reciprocal of the mean interarrival time (arrival rate). The fitting results presented in Table III show that the arrival rate at GHz is ns, which is higher than ns at GHz. It should be noted that these values are more than an order of magnitude larger than those observed in indoor corridor environments, owing to the outdoor setting.
In Fig. 11(b), the delay and power of the random components observed at each measurement location are normalized with respect to the delay and power of the LoS component at the same location. To model this characteristic, the path gain of the random component at delay and distance is as
where , represents the slope, and represents the intercept. Furthermore, and represent the delay and path gain of the LoS component at distance , respectively. denotes a zero-mean Gaussian random variable with variance , representing shadow fading. The parameters and are estimated by fitting the plot in Fig.11(b) using the maximum likelihood method [19], considering a noise level of dB. The results yield and at GHz, and and at GHz, illustrating a similar trend at both frequencies. Furthermore, as shown in Fig. 11(c), the path gain deviation to the model expressed in dB fits a normal distribution with a standard deviations of at GHz and at GHz, respectively. The obtained parameters for random cluster model are tabulated in Table III.
Fig. 12 presents an example of a CIR realization obtained through the proposed procedure, together with a comparison to the measured PDP in the delay domain (Rx3) at GHz. The reconstructed PDP, derived using the same signal bandwidth and incorporating the radiation patterns of the antennas employed in the measurement, shows close agreement with the measured PDP. Because the model incorporates a stochastic component, the two PDPs cannot coincide exactly; nevertheless, the results demonstrate the validity of the proposed model.
Scen. | Freq. | CI Model | FI Model |
LoS | GHz | ||
GHz | |||
NLoS | GHz | ||
GHz |
VI Path Loss and LSP Characterization
Regardless of the specific channel model, channel parameters including PL and LSPs are essential for applying standard channel models such as 3GPP TR 38.901 [2]. In this section, the measured channels are characterized in terms of PL, DS, AS, and Rician KF at both frequencies and in both scenarios.
VI-A Path Loss
The omnidirectional PL, is calculated as
where denotes the effect due to incremental angular steps below the HPBW and non-orthogonal radiation patterns of the used antennas which results in an overlap of the antenna beam which leads to an increase in the received power by as described in Recommendation ITU-R P.1407-8 [22]. In this study, was analytically calculated based on the measured antenna patterns. Note that the large measurement bandwidth allows almost all multipath components to be well separated in the delay domain. Consequently, the incoherent power summation in (14) yields a fading-free PL, irrespective of the antenna beamwidth. The model fitting parameters are evaluated in the following subsection.
To investigate the characteristics of the dual-band channels under different scenarios, two popular channel models: the close-in (CI) free-space reference distance model [20] and the floating intercept (FI) model [21], are used for data fitting.
CI model:
(15) where, is the Tx-Rx separation distance, is the path loss exponent (PLE), is the center frequency, and is the speed of light in .
FI model:
(16) where, and are the slope and floating intercept respectively.
The results of fitting the omnidirectional PLs for both LoS and NLoS scenarios using the aforementioned models are shown in Fig.13, and the corresponding model parameters are summarized in TableIV as well. From the fitting results, it can be observed that in the LoS scenario, both the CI and FI models yield similar parameter values, such as the PLE ( or ) and the standard deviation (), for each frequency band. In contrast, under the NLoS scenario, the values of the FI model for both the GHz and GHz bands are smaller than those of the CI model, suggesting that the FI model provides a better fit for NLoS conditions in this environment. Nonetheless, it is premature to draw definitive conclusions due to the limited number of samples.
From Fig. 13(a), it can be observed that the measured PL of each frequency band has distinctly exhibited distance dependency and several dB improvement can be achieved over FSPL due to multipath power in LoS scenario. Furthermore, when comparing the measured PL with the FSPL in each band, the difference at GHz is slightly smaller than that at GHz. This indicates that the GHz band undergoes somewhat greater degradation, which is consistent with the cluster characteristics discussed in the previous section.
In the NLoS scenario, the PLs are significantly higher, with a large shadowing distribution. This is evident from the standard deviation () values of the fitting parameters for both models, as shown in Fig. 13(b). At all positions, the PLs are considerably larger than the FSPL.
VI-B Large Scale Channel Parameters (LSPs)
The PDP and PAP are calculated using (1) and (3), respectively. The obtained power profiles are then processed according to an acceptance criterion. In particular, only those samples are retained that are either within dB of the peak sample in the power profiles [22] or within the cutoff level, whichever is higher. Similarly, the corresponding PAP is processed according to the same acceptance criterion before being used to evaluate the respective angular spreads (AS), namely the Azimuth-of-Departure (ASD) and the Azimuth-of-Arrival (ASA).
The temporal dispersion, on the other hand, is evaluated as the second moment of the PDP as follows
where each variable has the same meaning as has been explained earlier.
Similarly, ASD at Tx and ASA at Rx are obtained in terms of the relevant PAP as
where the representation of each variable is as per the earlier definition. In the above equations, if , then the corresponding refers to ASD and refers to ASA otherwise.
The power content in the LoS cluster compared to other significant MPCs is quantified as Rician KF and can be evaluated as
where represents the extracted complex path weight of the -th MPC within the -th cluster. denotes the LoS cluster.
The CDF plots of DS, ASD, ASA, and Rician KF for the considered scenarios are presented in Fig. 14. Based on the analysis, the key observations can be summarized as follows:
Due to several outlier MPCs with large delays observed at specific Rx positions, the DS exhibits large deviations above the 80% level; nevertheless, it remains similarly distributed at both frequencies up to this percentile.
The average ASA and ASD are approximately dB and dB, respectively, in the LoS scenario. They are similarly distributed at both frequencies; however, the ASA is consistently larger than the ASD, which can be attributed to the proximity of the Rx to the SW.
The K-factors range from to dB, with an average of to dB. No significant difference is observed between the distributions at GHz and GHz.
VII Conclusion
This paper presented an extensive D-D channel measurement campaign in a street-canyon environment at GHz and GHz, under both LoS and NLoS conditions, using an in-house-developed dual-band channel sounder. By applying clustering to merged datasets across the two frequency bands, both common and distinct multipath clusters were identified, enabling a comparative analysis of their frequency-dependent characteristics. The results show that the dominant scattering processes are primarily the direct path and single-bounce reflections from the north and south walls. Statistical analyses reveal no significant differences, but slightly greater degradation is observed at GHz.
A QD channel model was proposed, integrating deterministic components, LoS and major reflections from the north and south walls, with random components representing higher-order or irregular multipath contributions. Furthermore, LSPs, including PL, DS, AS, and the Rician KF, were characterized. These findings deepen the understanding of sub-THz propagation in urban street canyons and provide essential inputs for the development of accurate, frequency-dependent channel models, supporting the design and deployment of future 6G systems.